Company E decided to increase its share capital by capitalizing the goodwill of its fixed assets. To this end, the old shareholders will be given free of charge one new common stock and one new preferred share in each of the 2 old common and preferred shares respectively. The stock market price of the ordinary and preferred shares before the termination rightis is 45 euros and 30 euros respectively. Calculate the price of the ordinary and preference shares after deduction of the rights to receive and the value of the rights to receive each share.
Answer #1

Price per share After Issue of Additional 1 share for every 2 shares = Earlier Price per share*2/3

Price of Ordinary Share = 45*2/3 = 30 Euros per share

Price of Preference Share = 30*2/3 = 20 Euros per share